OpenStudy (anonymous):

1+2i multiplicity 2 does anyone know the additional zeros is it 1-2i multiplicity 2 or just 1-2i

OpenStudy (ikram002p):

1+2i multiplicity 2 means 2(1+2i) mm what is ur Qn exactly ?

OpenStudy (anonymous):

what the zeros are

OpenStudy (ikram002p):

zeros of what ? :D

OpenStudy (anonymous):

1+2i multiplicity 2

OpenStudy (ikram002p):

well 2(1+2i)=2+4i PS :- 1+2i is a complex number and i is not a variable \(i=\sqrt-1\)

OpenStudy (mathstudent55):

A polynomial with real coefficients that has complex roots has complex roots in complex conjugate pairs. If your polynomial has roots 1 + 2i multiplicity 2, that means it has the root 1 + 2i and 1 + 2i (the same root twice). Each of these two equal roots must come as a pair of complex conjugate roots, so there must also be the roots 1 - 2i and 1 - 2i, or 1 - 2i multiplicity 2.

OpenStudy (ikram002p):

mmm is that was her Qn !! lolz like a buzzle